a cylindrical capacitor has a length of 6cm is made of two concentric rings with an inner radius of 2.5 cm and an outer radius of 3.5cm. how much charge is this capacitor id it is connected to a 15V battery?


1
Expert's answer
2021-04-26T17:04:54-0400

The capacitance of cylindrical capacitor is "C = \frac{2 \pi \varepsilon \varepsilon_0}{\ln \frac{b}{a}} L".

From the definition of capacitance, the charge on each plate of the capacitor is: "q = C U = \frac{2 \pi \varepsilon \varepsilon_0}{\ln \frac{b}{a}} L U = \frac{2 \cdot 3.14 \cdot 8.85 \cdot 10^{-12}\frac{F}{m}}{\ln \frac{3.5}{2.5}} \cdot 0.06 m \cdot 15 V \approx 148.7 \cdot 10^{-12} C = 148.7 \ pC".